The Collins Group, a leading producer of custom automobile accessories, has hired you to estimate the firm's weighted average cost of capital. The balance sheet and some other information are provided below.

  Assets
Current assets $ 38,000,000
Net plant, property, and equipment 101,000,000
Total assets $139,000,000
  Liabilities and Equity
Accounts payable $ 10,000,000
Accruals 9,000,000
Current liabilities $ 19,000,000
Long-term debt (40,000 bonds, $1,000 par value) 40,000,000
Total liabilities $ 59,000,000
Common stock (10,000,000 shares) 30,000,000
Retained earnings 50,000,000
Total shareholders' equity 80,000,000
Total liabilities and shareholders' equity $139,000,000


The stock is currently selling for $15.25 per share, and its noncallable $1,000 par value, 20-year, 7.25% bonds with semiannual payments are selling for $875.00. The beta is 1.25, the yield on a 6-month Treasury bill is 3.50%, and the yield on a 20-year Treasury bond is 5.50%. The required return on the stock market is 11.50%, but the market has had an average annual return of 14.50% during the past 5 years. The firm's tax rate is 40%.

Refer to the data for the Collins Group. Based on the CAPM, what is the firm's cost of common stock?

a. 13.65%
b. 11.15%
c. 13.00%
d. 12.35%
e. 11.73%
Answer #1

Given :

Risk Free Rate (Rf) = 5.50%

Return on Stock Market (Rm) = 11.50%

Beta (B) = 1.25

Cost of Equity = Rf + (Rm - Rf)B

= 5.50% + (11.50%-5.50%)1.25

= 13%

Cost of Debt :

Let

  • r = semiannual interest rate
  • P = bond price = $875
  • C = bond coupon = 7.25% / 2 * 1,000 = $36.25
  • n = semiannual number of periods = 20 * 2 = 40

The semiannual yield can be found with the help of the following bond price formula:

Bond Price = Bond Coupon * [1-(1+YTM)-n]/YTM + Face Value/(1+YTM)n

YTM = 4.28%

The bond's yield to maturity (YTM) is the semiannual yield multiplied by two:

  • YTM = 2 * r
  • YTM = 2 * 4.28%
  • YTM = 8.56%

Therefore, the after-tax cost of debt is:

  • After-tax cost of debt = YTM * (1 - Tax rate)
  • After-tax cost of debt = 8.56% * (1 - 0.4)
  • After-tax cost of debt = 5.14%

The after-tax cost of debt is 5.14%

The WACC = 40mn/120mn * 5.14% + 80m/120m * 13%

= 10.38%